# The Procrustean Bed of Time Series: The Optimization Bias of Point-wise Loss

### v1

@@ G.1 Characteristic of Discrete Fourier Transform
$\frac{\partial f}{\partial x}=U^{H}.$ (90)
The DFT is uniquely suited to our framework for three key advantages that directly align with the goals of proposed debiasing program:
@@ G.2 Characteristic of Discrete Wavelet Transform
As an orthogonal operation, its inverse is its transpose ($W^{T}$), which facilitates seamless backpropagation during training:
$\frac{\partial w}{\partial x}=W^{T}$ (92)
@@ G.4.1 Point Estimation Unbiasedness
$\mathcal{L}_{\text{Harm},\ell_{1}}^{k}=w_{k}\|f_{k}-\hat{f}_{k}\|_{2},\quad\text{where}\quad w_{k}=1+\bar{f}_{k}$ (99)
@@ G.4.2 Element Bias: Balancing Training Stability and Dynamics Rectification
We generalize Spectral Bias to Element Bias to describe the phenomenon where optimization dynamics are dominated by high-magnitude components in any structural orthogonalization (e.g., DFT, DWT, or SVD). While a strictly unbiased loss (i.e., spectral whitening, $\mathcal{L}_{\text{Unbiased},\ell_{2}}=\sum_{k}^{K}\|f_{k}-\hat{f}_{k}\|_{2}^{2}/\bar{f}_{k}^{2}$ and $\mathcal{L}_{\text{Unbiased},\ell_{1}}=\sum_{k}^{K}\|f_{k}-\hat{f}_{k}\|_{2}/\bar{f}_{k}$) theoretically ensures uniform gradient sensitivity, it is practically pathological due to the uneven Signal-to-Noise Ratio (SNR) distribution. Specifically, strict whitening assigns aggressive weights to noise-dominated low-amplitude components, leading to training instability (gradient explosion) and noise overfitting, while simultaneously causing the optimizer to underfit the crucial high-amplitude main structures due to their relatively lower gradient contribution.
@@
  • Stability Anchor (Dominant Regime): For large-amplitude components, the amplitude ($\nabla_{k}\mathcal{L}_{\text{Harm},\ell_{2}}\propto\bar{f}_{k}$, $\nabla_{k}\mathcal{L}_{\text{Harm},\ell_{1}}\propto\gamma\bar{f}_{k}$) harmonized gradient. This deliberately retains the standard norm’s focus on fundamental structures, anchoring the optimization trajectory and preventing destabilization by high-frequency noise.
  • Dynamics Rectification (Weak Regime): For small-amplitude components, the harmonized gradient converges to unity ($\nabla_{k}\mathcal{L}_{\text{Harm},\ell_{2}}\approx\gamma$, $\nabla_{k}\mathcal{L}_{\text{Harm},\ell_{1}}\approx 1$). This acts as an unbiased correction to rectify gradient fatigue, ensuring that subtle but informative details receive sufficient gradient flow to be learned effectively.
@@ H.1 Point-wise loss function based on $\ell_{2}$ Norm
H.1.1 Independent Optimization of Real and Imaginary Components
For a time series sample $x\in\mathbb{R}^{L}$. DFT can be expressed as a linear projection via a unitary matrix $U\in\mathbb{C}^{K\times L}$:
@@ H.1.1 Independent Optimization of Real and Imaginary Components
where $f\in\mathbb{C}^{K}$ is the complex spectrum. The elements of the DFT matrix $U$ are given by $u_{k,\tau}=\frac{1}{\sqrt{T}}\exp(-j2\pi k\tau/T)$. Notably, $U$ satisfies $U^{H}U=\mathbf{I}$, where $U^{H}$ is the conjugate transpose of $U$.
We can decompose $U$ into real and imaginary parts, $U=U^{r}+jU^{i}$. A common strategy for frequency-domain optimization is to separately penalize the errors in the real and imaginary components of the spectrum:
@@
where $\hat{x}$ is the model’s predicted and $\hat{f}=U\hat{x}$.
However, a closer examination of this objective reveals a critical flaw. By leveraging the properties of complex numbers and the unitary nature of $U$, we can simplify $\mathcal{L}_{\text{freq}}$:
$\mathcal{L}_{\text{freq},\ell_{2}}$ $=\|U^{r}(x-\hat{x})\|_{2}^{2}+\|U^{i}(x-\hat{x})\|_{2}^{2}$ (105)
$=\|U(x-\hat{x})\|_{2}^{2}=\|x-\hat{x}\|_{2}^{2}$
$\equiv\mathcal{L}_{\text{temp},\ell_{2}}.$
This derivation shows that minimizing the squared $\ell_{2}$ error in the frequency domain is equivalent to minimizing it in the time domain. Consequently, their gradients are identical:
@@
This results also can be explained via Parseval’s theorem.
@@
(Ineffectiveness of Unitary Transforms for $\ell_{2}$ Orthogonality) For any unitary transform, minimizing the squared $\ell_{2}$ norm of the transformed error is mathematically equivalent to minimizing the squared $\ell_{2}$ norm of the original error. Therefore, such transforms cannot eliminate optimization bias.
H.1.2 Independent Optimization of Amplitude and Phase
An alternative approach is to disentangle the objectives by penalizing errors in the amplitude and phase spectra separately. Let the predicted amplitude and phase be $\hat{A}$ and $\hat{\theta}$, respectively,
$A=|Ux|,\quad\theta=\arg(Ux),$ (107)
@@ H.1.2 Independent Optimization of Amplitude and Phase
The gradients of these loss functions with respect to the prediction $\hat{x}$ are:
[…]
$\frac{\partial\mathcal{L}_{\theta,\ell_{2}}}{\partial\hat{x}}$ $=+2(U^{i})^{\top}\left((\theta-\hat{\theta})\odot\frac{U^{r}\hat{x}}{\hat{A}^{2}}\right)-2(U^{r})^{\top}\left((\theta-\hat{\theta})\odot\frac{U^{i}\hat{x}}{\hat{A}^{2}}\right)$ (110)
where $\odot$ denotes the Hadamard product.
H.1.3 Independent Optimization of Error Amplitude and Phase
We now consider the amplitude and phase of the error signal itself, $e=x-\hat{x}$:
$\bar{A}=|Ue|,\quad\bar{\theta}=\arg(Ue)$ (111)
Analyzing the gradient of the error amplitude loss, $\mathcal{L}_{\bar{A}}$, reveals another equivalence:
$\mathcal{L}_{\bar{A},\ell_{2}}=\|U(x-\hat{x})\|_{2}^{2}=\|x-\hat{x}\|_{2}^{2}=\mathcal{L}_{\text{temp}}$ (113)
[…]
Once again, simply optimizing the amplitude of the frequency-domain error under the $\ell_{2}$ norm collapses back to the original temporal MSE objective.
However, the gradient for the phase of the error, $\mathcal{L}_{\bar{\theta}}$, provides a unique learning signal:
This demonstrates that while penalizing the error amplitude ($\mathcal{L}_{\bar{A}}$) is redundant, penalizing the error phase ($\mathcal{L}_{\bar{\theta}}$) offers a genuinely novel objective.
H.2 Point-wise loss function based on $\ell_{1}$ Norm
We now investigate whether using the $\ell_{1}$ norm, known for its robustness and different gradient characteristics, can yield more effective disentanglement. The subgradient of the $\ell_{1}$ norm is the sign function, which provides updates of constant magnitude, unlike the $\ell_{2}$ norm’s error-proportional updates.
H.2.1 Independent Optimization of Real and Imaginary Components
The $\ell_{1}$ analogue to the first $\ell_{2}$ loss is:
$\mathcal{L}_{\text{freq},\ell_{1}}=\|U^{r}(x-\hat{x})\|_{1}+\|U^{i}(x-\hat{x})\|_{1}.$ (116)
Its subgradient with respect to $\hat{x}$ is:
where $e=x-\hat{x}$. This gradient does not collapse to the temporal $\ell_{1}$, indicating that it provides a distinct optimization path.
H.2.2 Independent Optimization of Amplitude and Phase
Applying the $\ell_{1}$ norm to the amplitude and phase of the predictions yields the following objectives:
The corresponding subgradients are:
@@ H.2.2 Independent Optimization of Amplitude and Phase
H.2.3 Independent Optimization of Error Amplitude and Phase
Finally, we consider the $\ell_{1}$ norm of the error’s amplitude and phase:
@@ H.2.3 Independent Optimization of Error Amplitude and Phase
The subgradients are:
$\frac{\partial\mathcal{L}_{\bar{A},\ell_{1}}}{\partial\hat{x}}$ $=-(U^{r})^{\top}\left(\operatorname{sgn}(\bar{A})\odot\frac{U^{r}e}{\bar{A}}\right)-(U^{i})^{\top}\left(\operatorname{sgn}(\bar{A})\odot\frac{U^{i}e}{\bar{A}}\right)$ (122)
$=-(U^{r})^{\top}\left(\delta(\bar{A})\odot\frac{U^{r}e}{\bar{A}}\right)-(U^{i})^{\top}\left(\delta(\bar{A})\odot\frac{U^{i}e}{\bar{A}}\right)-\frac{e}{\bar{A}}$
$\approx-\frac{e}{\bar{A}}$
where $\delta(x)=0$ is the Kronecker Delta function, defined as $\delta(x)=\begin{cases}1,&\text{if}\;x=0\\ 0,&\text{otherwise}\end{cases}$. The approximate result assumes $\bar{A}>0$ for all $k$, the gradient is a vector containing only phase information.
@@
Both gradients provide non-trivial and distinct learning signals compared to a simple temporal $\ell_{1}$ loss, confirming that the $\ell_{1}$ norm offers more robust options for objective disentanglement in the frequency domain.
@@ I.1.2 Statistical Property of Stochastic Component
$\textit{SSNR}_{z}=\frac{\sigma_{z}^{2}}{\sigma_{\epsilon}^{2}}=\frac{1}{1-\phi_{i}^{2}}\Rightarrow\phi_{1}=\sqrt{\frac{\textit{SSNR}_{z}-1}{\textit{SSNR}_{z}}}.$ (134)
@@
$\textit{SSNR}_{x}=\textit{SSNR}_{v}+\textit{SSNR}_{z}=\frac{KA^{2}}{2\sigma_{\epsilon}^{2}}+\frac{1}{1-\phi^{2}}.$ (135)
@@ I.4.1 For EOB Theory Verification Experiment
Poisson Approx. Sym. $\{0,1,\dots\}$ Middle $\approx 3$ $\lambda=0.5$
@@
I.4.2 For Long-term Forecasting Tasks

### v2

@@ G.1 Characteristic of Discrete Fourier Transform
$\frac{\partial f}{\partial x}=U^{\mathrm{H}}.$ (88)
The DFT is uniquely suited to our framework for two key advantages that directly align with the goals of proposed debiasing program:
@@ G.2 Characteristic of Discrete Wavelet Transform
As an orthogonal operation, its inverse is its transpose ($W^{\top}$), which facilitates seamless backpropagation during training:
$\frac{\partial w}{\partial x}=W^{\top}$ (90)
@@ G.4.1 Point Estimation Unbiasedness
$\mathcal{L}_{\text{Harm},\ell_{1}}^{k}=w_{k}\|f_{k}-\hat{f}_{k}\|_{2},\quad\text{where}\quad w_{k}=1+\gamma\bar{f}_{k}$ (97)
@@ G.4.2 Element Bias: Balancing Training Stability and Dynamics Rectification
We generalize Spectral Bias to Element Bias to describe the phenomenon where optimization dynamics are dominated by high-magnitude components in any structural orthogonalization (e.g., DFT, DWT, or SVD). While a strictly unbiased loss (i.e., spectral whitening, $\mathcal{L}_{\text{Unbiased},\ell_{2}}=\sum_{k}^{K}\|f_{k}-\hat{f}_{k}\|_{2}^{2}/\bar{f}_{k}^{2}$ and $\mathcal{L}_{\text{Unbiased},\ell_{1}}=\sum_{k}^{K}\|f_{k}-\hat{f}_{k}\|_{1}/\bar{f}_{k}$) theoretically ensures uniform gradient sensitivity, it is practically pathological due to the uneven Signal-to-Noise Ratio (SNR) distribution. Specifically, strict whitening assigns aggressive weights to noise-dominated low-amplitude components, leading to training instability (gradient explosion) and noise overfitting, while simultaneously causing the optimizer to underfit the crucial high-amplitude main structures due to their relatively lower gradient contribution.
@@
  • Stability Anchor (Dominant Regime): For large-amplitude components where $\bar{f}_{k}\gg\gamma$, the harmonizing weight approaches unity ($w_{k}\to 1$). Consequently, the gradient behavior converges to that of the standard norm ($\nabla_{k}\mathcal{L}_{\text{Harm},\ell_{2}}\propto(f_{k}-\hat{f}_{k})=e_{k}\bar{f}_{k}$, $\mathcal{L}_{\text{Harm},\ell_{1}}\propto\gamma\bar{f}_{k}$), retaining sensitivity to amplitude. This regime deliberately anchors the optimization trajectory on fundamental high-energy structures, preventing destabilization by high-frequency noise.
  • Dynamics Rectification (Weak Regime): For small-amplitude components where $\bar{f}_{k}\ll\gamma$, the harmonized gradient converges to unity ($\nabla_{k}\mathcal{L}_{\text{Harm},\ell_{2}}\approx\gamma$, $\nabla_{k}\mathcal{L}_{\text{Harm},\ell_{1}}\approx 1$). This acts as an unbiased correction to rectify gradient fatigue, ensuring that subtle but informative details receive sufficient gradient flow regardless of their scale.
@@ H.1 Point-wise loss function based on $\ell_{2}$ Norm
H.1.1 Decoupled Optimization of Real and Imaginary Components
Consider a time series sample $x\in\mathbb{R}^{T}$. DFT can be expressed as a linear projection via a unitary matrix $U\in\mathbb{C}^{T\times T}$:
@@ H.1.1 Decoupled Optimization of Real and Imaginary Components
where $f\in\mathbb{C}^{T}$ is the complex spectrum. The elements of the DFT matrix $U$ are given by $u_{k,\tau}=\frac{1}{\sqrt{T}}\exp(-j2\pi k\tau/T)$. Notably, $U$ satisfies the unitary property $U^{\mathrm{H}}U=\mathbf{I}$, where $U^{\mathrm{H}}$ denotes the conjugate transpose.
We can decompose the spectrum into real and imaginary parts, $f=f^{r}+jf^{i}$. A common intuitive strategy for frequency-domain optimization is to separately penalize the errors in these components:
@@
where $\hat{x}$ is the model’s prediction and $\hat{f}=U\hat{x}$.
However, a closer examination reveals that this objective is mathematically redundant under the $\ell_{2}$ norm. Leveraging the property that for any complex vector $z$, $\|z\|_{2}^{2}=\|\Re(z)\|_{2}^{2}+\|\Im(z)\|_{2}^{2}$, and the unitary nature of $U$, we can simplify $\mathcal{L}_{\text{freq}}$:
$\mathcal{L}_{\text{freq},\ell_{2}}$ $=\|f-\hat{f}\|_{2}^{2}$ (103)
$)\displaystyle=(U(x-\hat{x}))^{\mathrm{H}}(U(x-\hat{x}$))
@@
$=\|x-\hat{x}\|_{2}^{2}\equiv\mathcal{L}_{\text{temp},\ell_{2}}.$
This derivation proves that minimizing the squared $\ell_{2}$ error in the frequency domain is exactly equivalent to minimizing it in the time domain. Consequently, their gradients are identical:
[…]
This result is a direct manifestation of Parseval’s Theorem, which guarantees the preservation of energy (and thus Euclidean distance) under unitary transformations.
@@
(Invariance of $\ell_{2}$ Optimization under Unitary Transform) Let $U$ be a unitary operator (e.g., DFT matrix) such that $U^{\mathrm{H}}U=I$. The optimization bias under the squared $\ell_{2}$ norm (MSE) is invariant to the transformation $U$.
Remark: (Reconciliation with Strategy 3) Theorem H.1 demonstrates that simply changing the basis does not reduce optimization bias if the loss function remains the standard MSE ($\ell_{2}$ norm), as the gradient dynamics are merely rotated and isometric. However, the proposed debiasing method in Section 4 combines this transform with non-isotropic loss functions. In the transformed domain where components are orthogonal, a component-wise loss function acts on the decoupled modes, thereby breaking the rotational invariance and altering the effective SSNR experienced by the optimizer.
H.1.2 Decoupled Optimization of Amplitude and Phase
[…]
@@ H.1.2 Decoupled Optimization of Amplitude and Phase
The gradients of these loss functions with respect to the prediction $\hat{x}$ are derived via the chain rule. We can simplify the expressions by noting that $U^{r}\hat{x}/\hat{A}=\cos\hat{\theta}$ and $U^{i}\hat{x}/\hat{A}=\sin\hat{\theta}$:
@@
where $\odot$ denotes the Hadamard product.
Remark: The term $\hat{A}$ in the denominator of the phase gradient reveals a critical numerical instability: for frequency components with near-zero amplitude (low energy), the phase gradient tends to explode.
H.1.3 Decoupled Optimization of Error Amplitude and Phase
We now examine the properties of the error signal itself in the frequency domain. Let the error vector be $e=x-\hat{x}$, and its complex spectrum be $f_{e}=Ue$. We define the error amplitude $\bar{A}\in\mathbb{R}^{T}$ and error phase $\bar{\theta}\in\mathbb{R}^{T}$ as:
$\bar{A}=|f_{e}|=|U(x-\hat{x})|,\quad\bar{\theta}=\arg(f_{e}).$ (108)
The corresponding squared $\ell_{2}$ objectives are:
1. The Redundancy of Error Amplitude: Analyzing the gradient of the error amplitude loss reveals a fundamental equivalence. By Parseval’s theorem, the energy of the error spectrum magnitude is identical to the energy of the time-domain error:
$\mathcal{L}_{\bar{A},\ell_{2}}=\|U(x-\hat{x})\|_{2}^{2}=\|x-\hat{x}\|_{2}^{2}\equiv\mathcal{L}_{\text{temp},\ell_{2}}.$ (110)
Consequently, its gradient collapses back to the standard residual:
[…]
Insight: Simply optimizing the amplitude of the frequency-domain error under the $\ell_{2}$ norm yields no structural advantage, as it is mathematically isomorphic to the original temporal MSE objective.
2. The Distinctness of Error Phase: In contrast, the gradient for the phase of the error, $\mathcal{L}_{\bar{\theta}}$, provides a mathematically distinct learning signal:
This confirms that while penalizing error amplitude is redundant, penalizing error phase offers a genuinely novel optimization path that is orthogonal to standard MSE.
Remark (Gradient Instability): Note the term $\bar{A}^{2}$ in the denominator. As the model converges and the error amplitude $\bar{A}\to 0$, the gradient for the phase objective approaches infinity. This singularity suggests that pure phase optimization is numerically unstable near convergence.
H.2 Point-wise loss function based on $\ell_{1}$ Norm
We now investigate the properties of the $\ell_{1}$ norm in the frequency domain. Unlike the $\ell_{2}$ norm, which induces isotropic gradients leading to the time-frequency equivalence proven in Theorem H.1, the $\ell_{1}$ norm is anisotropic. Its subgradient (i.e., the sign function) provides update signals that are independent of the error magnitude, offering distinct robustness and sparsity-inducing characteristics.
H.2.1 Decoupled Optimization of Real and Imaginary Components
The $\ell_{1}$ analogue to the standard spectral loss is defined by separately penalizing the real and imaginary components:
$\mathcal{L}_{\text{freq},\ell_{1}}=\|\Re(f-\hat{f})\|_{1}+\|\Im(f-\hat{f})\|_{1}=\|U^{r}(x-\hat{x})\|_{1}+\|U^{i}(x-\hat{x})\|_{1}.$ (113)
Its subgradient with respect to the temporal prediction $\hat{x}$ is derived as:
where $e=x-\hat{x}$ is the residual vector.
The Breaking of Equivalence: Crucially, unlike the $\ell_{2}$ case where Parseval’s theorem guarantees $\|Ux\|_{2}=\|x\|_{2}$, the $\ell_{1}$ norm does not satisfy rotational invariance (i.e., $\|Ux\|_{1}\neq\|x\|_{1}$). Geometrically, the $\ell_{1}$ unit ball is a diamond shape, which changes orientation relative to the axes under rotation $U$. Consequently, the gradient in Eq. (113) does not collapse to the temporal $\ell_{1}$ update. This confirms that frequency-domain $\ell_{1}$ optimization provides a mathematically distinct trajectory that effectively exploits the sparsity of spectral components.
H.2.2 Decoupled Optimization of Amplitude and Phase
[…]
The corresponding subgradients are derived via the chain rule. We can simplify the expressions by noting that $U^{r}\hat{x}/\hat{A}=\cos\hat{\theta}$ and $U^{i}\hat{x}/\hat{A}=\sin\hat{\theta}$:
[…]
Geometric Interpretation:
  • Amplitude Gradient: The update vector is directed along the current phase angle $\hat{\theta}$ with a constant magnitude (determined by $\operatorname{sgn}$). This pushes the signal energy towards the target without being diminished by the gradient dominance of large peaks, a common issue in $\ell_{2}$.
  • Phase Gradient: The update is orthogonal to the amplitude direction. Crucially, it scales with $1/\hat{A}$.
Remark: The presence of $1/\hat{A}$ in the phase gradient again highlights the Phase Singularity problem: for frequency components where the model predicts near-zero energy ($\hat{A}\approx 0$), the phase gradient becomes numerically unstable. This confirms that while $\ell_{1}$ solves the scale sensitivity issue, it does not resolve the small-signal singularity.
H.2.3 Decoupled Optimization of Error Amplitude and Phase
Finally, we consider the $\ell_{1}$ norm of the error’s amplitude and phase. Let $\bar{A}=|Ue|$ be the amplitude of the error spectrum.
@@ H.2.3 Decoupled Optimization of Error Amplitude and Phase
Gradient of Error Amplitude: Since the error amplitude $\bar{A}$ is intrinsically non-negative, minimizing its $\ell_{1}$ norm is equivalent to minimizing the sum of magnitudes. The subgradient derivation simplifies significantly compared to $\ell_{2}$ case:
$\frac{\partial\mathcal{L}_{\bar{A},\ell_{1}}}{\partial\hat{x}}$ $=-(U^{r})^{\top}\left(\frac{\partial\bar{A}}{\partial\bar{A}}\odot\frac{U^{r}e}{\bar{A}}\right)-(U^{i})^{\top}\left(\frac{\partial\bar{A}}{\partial\bar{A}}\odot\frac{U^{i}e}{\bar{A}}\right)$ (119)
$=-(U^{r})^{\top}\left(\textbf{1}\odot\cos\bar{\theta}\right)-(U^{i})^{\top}\left(\textbf{1}\odot\sin\bar{\theta}\right)$
$=-\text{IDFT}\left(\frac{Ue}{|Ue|}\right)=-\text{IDFT}(e^{j\bar{\theta}})$
where $\bar{\theta}$ is the phase of the error vector.
Interpretation: This result is theoretically profound. The gradient depends solely on the phase of the error, completely ignoring the magnitude scale. This effectively acts as a “Spectral Whitener” during backpropagation, applying uniform gradient pressure across all frequencies regardless of their energy. This prevents high-energy low-frequency components from dominating the optimization, a key advantage over $\ell_{2}$ based losses.
Gradient of Error Phase: Similarly, the gradient for the error phase focuses purely on rotational alignment:
@@
Both gradients provide non-trivial and distinct learning signals compared to a simple temporal $\ell_{1}$ loss, confirming that the frequency-domain $\ell_{1}$ norm offers robust mechanisms for objective disentanglement.
@@ I.1.2 Statistical Property of Stochastic Component
$\textit{SSNR}_{z}=\frac{\sigma_{z}^{2}}{\sigma_{\epsilon}^{2}}=\frac{1}{1-\phi_{1}^{2}}\Rightarrow\phi_{1}=\sqrt{\frac{\textit{SSNR}_{z}-1}{\textit{SSNR}_{z}}}.$ (131)
@@
$\textit{SSNR}_{x}=\textit{SSNR}_{v}+\textit{SSNR}_{z}=\frac{KA^{2}}{2\sigma_{\epsilon}^{2}}+\frac{1}{1-\phi_{1}^{2}}.$ (132)
@@ I.4.1 For EOB Theory Verification Experiment
As described in Eq. (123), we fix the stochastic component $z_{t}$ (constant $\sigma_{z}^{2}$) and increase the amplitude $A$ of the deterministic component $v_{t}$. While this effectively increases SSNR, it also increases the total variance $\sigma_{x}^{2}$. This protocol highlights the model’s inability to converge to the theoretical optimum $\text{MSE}^{\text{opt.}}$ as the deterministic signal grows.
@@
Poisson Approx. Sym. $\{0,1,\dots\}$ Middle $\approx 3$ $\lambda=0.25$
@@
I.4.2 For Insight Experiments
The generated trigonometric series follows the dynamic mechanism described in Appendix I.1.1. We randomly select $k\in\{1,2,3,4,5\}$ components with a maximum cutoff frequency of $f_{\max}=15$ and a random phase $\theta$. The original series is split into training and testing sets with a ratio 7:3. Both the historical context and the forecasting horizon are set to 200 points. A Transformer architecture is employed as the model backbone.
I.4.3 For Long-term Forecasting Tasks